If the number of flaws on a piece of plywood occurs randomly and follows the Poisson distribution with a mean of 1 per 5m^2

.

(a) Find the probability that a piece of randomly chosen plywood will contain no flaws.

(b) The plywood is sold in pieces of size 3m^2. Find the probability that the plywood will contain no flaws.


1
Expert's answer
2021-03-18T13:58:17-0400

"\text{Poisson's formula:}"

"P(X=m)=\frac{\lambda^m{e^{-\lambda}}}{m!}"

"a)\text{ lets a piece of plywood has an area } S"

"\text{then }\lambda=\frac{S}{5}"

"P(0) -\text{the probability that the plywood does not contain flaws}"

"P(0)=\frac{\lambda^0{e^{-\lambda}}}{0!}=e^{-\lambda}=e^{-\frac{S}{5}}(1)"

"b)\text{we use the formula (1)}"

"P(0)= e^{-\frac{3}{6}}\approx0.61"

Answer:

"a)P(0)=e^{-\frac{S}{5}},\text{where S area of a piece of plywood}\newline
b)P(0)=0.6"